OpenStudy (anonymous):

When using the substitution method, what expression for x in the second equation would be substituted in the first equation? 4x – 7y = –9 x + 3y = 12 –3y + 12 3y – 12 –3y – 12 3y + 12

OpenStudy (amistre64):

solve for x in the second equation ...

OpenStudy (amistre64):

i think in chapter 3 or 4 they call it, "make x the subject"

OpenStudy (anonymous):

but how will that get me an operation?

OpenStudy (noseboy908):

-3y+12. to get x alone on one side of the second eqn., subtract 3y from each side to get -3y+12=x.

OpenStudy (amistre64):

by substituting a "value" for x into the first equation that is modeled after the second equation, we keep the overall value of the system the same; but the substitution from the second into the first will eliminate the x variable and leave us with only y variables to solve.

OpenStudy (amistre64):

4x – 7y = –9 x + 3y = 12 --> x = 12-3y 4x – 7y = –9 4(12-3y) – 7y = –9 48 -12y – 7y = –9 48 -19y = –9 etc ....